OpenStudy (anonymous):

subtract and simplify 3√45-√20-√80 i have it broken down to (3)3√9√5-2√2√5-√14√5 i know the answer is 3√5 but don't know how to get to 3√5

OpenStudy (anonymous):

It breaks down to: \[9\sqrt{5} - 2\sqrt{5} - 4\sqrt{5} = 3\sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{45}=\sqrt{3*3*5}=3\sqrt{5}\]\[\sqrt{20}=\sqrt{2*2*5}=2\sqrt{5}\]\[\sqrt{80}=\sqrt{4*4*5}=4\sqrt{5}\]so you subtraction becomes \[3\sqrt{5}-2\sqrt{5}-4\sqrt{5}=-3\sqrt{5}\]

OpenStudy (anonymous):

sorry there was one 3 with sqrt(45) so ans will be

OpenStudy (anonymous):

\[+3\sqrt{5}\]

OpenStudy (anonymous):

ahhh i see now i calculated wrong for √80 and that was what was throwing me off.